The difference of two numbers is 91. The second is 5 less than 5 times the first. What are the two numbers?
Answer by checkley71(8403)

X-Y=91 OR X=91+Y
Y=5X-5 NOW SUBSTITUTE (91+Y) FOR X
Y=5(91+Y)-5
Y=455+5Y-5
Y-5Y=-5-455
-4Y=-460
Y=-460/-4
Y=115 ANSWER.
115=5X-5
5X=115+5
5X=120
X=120/5
X=24 ANSWER.
115-24=91
91=91